OpenStudy (anonymous):

How do you find the integral of 4/(x^2+2x+5) using trig substitution?

OpenStudy (anonymous):

start by completing the squares for the denominator

OpenStudy (anonymous):

okay; that makes sense now. thanks!

OpenStudy (anonymous):

anymore questions?

OpenStudy (anonymous):

no :) I just didn't know how to start the problem.

sam (.sam.):

\[4 \int\limits \frac{1}{x^2+2 x+5} \, dx\] CTS \[4 \int\limits \frac{1}{4+(x+1)^2} \, dx\] Factor 4 \[ \int\limits \frac{1}{ \left(1+\frac{u^2}{4}\right)} \, du\] Trig substitution here

OpenStudy (anonymous):

ok, thanks. so it would be 2arctan((x+1)/2) + c

OpenStudy (anonymous):

Yes.